# Spurious reconstruction from brain activity

### v3

@@ Abstract
Advances in brain decoding, particularly visual image reconstruction, have sparked discussions about the societal implications and ethical considerations of neurotechnology. As these methods aim to recover visual experiences from brain activity and achieve prediction beyond training samples (zero-shot prediction), it is crucial to assess their capabilities and limitations to inform public expectations and regulations. Our case study of recent text-guided reconstruction methods, which leverage a large-scale dataset (Natural Scene Dataset, NSD) and text-to-image diffusion models, reveals limitations in their generalizability. We found poor performance when applying these methods to a different dataset designed to prevent category overlaps between training and test sets. UMAP visualization of the text features with NSD images showed a limited diversity of semantic and visual clusters, with overlap between training and test sets. Formal analysis and simulations demonstrated that clustered training samples can lead to “output dimension collapse,” restricting predictable output feature dimensions. Simulations further showed that diversifying the training set improved generalizability. However, text features alone are insufficient for mapping to the visual space. We argue that recent realistic reconstructions may primarily be a blend of classification into trained categories and generation of inauthentic images through text-to-image diffusion (hallucination). Diverse datasets and compositional representations spanning the image space are essential for genuine zero-shot prediction. Interdisciplinary discussions grounded in understanding the current capabilities and limitations, as well as ethical considerations, of the technology are crucial for its responsible development.
Author summary
Brain decoding has gained attention not only in neuroscience but also in fields like computer science, and even policy-making. However, claims about its predictive performance often tend to be overly optimistic. Here, we critically examine recent brain decoding methods, particularly visual image reconstruction methods that use generative AIs to generate seemingly realistic images from brain activity. Our analyses reveal that the apparent success of these methods largely stems from exploiting similarities between training and test stimuli in the experiment and from the hallucination of recent generative AIs, failing to achieve the genuine zero-shot reconstruction of novel images. We introduce the concept of “output dimension collapse,” a fundamental limitation in predicting within a target space, and discuss conditions necessary for true zero-shot prediction in terms of compositional representations and image recovery from latent features. We also highlight the limitations of current evaluation metrics and provide a framework for interpreting seemingly realistic reconstructions. By examining current limitations and proposing improvement strategies, we aim to encourage rigorous development and evaluation of brain decoding techniques.
Introduction
@@
Brain decoding has been widely used in the neuroscience field, revealing specific contents of the mind [1, 2, 3, 4]. As brain decoding is sometimes referred to as “mind-reading” in popular media [5, 6, 7], it has attracted significant attention beyond the scientific community due to its potential for real-world applications in medicine and industry. Such neurotechnology has also started to affect future ethical discussions and legal regulations [8]. To prevent misleading public expectations and policies, scientists need to carefully assess the current status of brain decoding techniques and clarify the possibilities and limitations.
@@
One of the major challenges in brain decoding is the limited amount of brain data we can collect. The current brain measurement devices are costly, yielding far less brain data than the amounts typically used in image or text processing within the field of computer science and AI [9, 10]. Although we have gradually increased the amount of brain data per subject [11, 12, 13, 14, 15], it remains impractical to collect brain data covering the full range of cognitive states and perceptual experiences. Consequently, classification-based decoding approaches, primarily developed in the early stages of this field, are insufficient to uncover the neural code in general or natural conditions since decodable information is confined to the same stimuli or predefined categories used in a training phase.
@@
To overcome this limitation, several decoding methods have been developed to enable the prediction of novel contents from brain activities that are not encountered during the training phase. Kay et al. [16] proposed a general visual decoding approach via a statistical encoding model that predicted fMRI voxel values from image features. It successfully identified novel test images from a set of 1,000 candidates. Mitchell et al. [17] utilized co-occurrence rates of specific verb sets for nouns and built a computational model to predict fMRI voxel values while thinking about nouns in presented line drawing images. Their model demonstrated the ability to predict voxel values for novel nouns not seen during the training phase. Brouwer and Heeger [18] constructed a color-tuning model and predicted brain activity while the subjects were presented with color stimuli. As their training stimuli covered most of the color space, their methods successfully identified novel colors not included in the training dataset. Horikawa and Kamitani [19] utilized deep neural network (DNN) features to decode brain activity measured while subjects perceived natural images. They showed successful prediction of novel object categories not encountered during the training phase.
@@
In the field of machine learning, “zero-shot” prediction refers to the ability of a model to accurately predict or classify novel contents not encountered during the training phase [20, 21]. This capability has emerged in various applications across different domains, including image classification [22], image generation [23], and natural language processing [24]. The concept of zero-shot prediction can be considered analogous to brain decoding techniques that aim to interpret brain activity patterns associated with previously unseen stimuli or experiences. Both approaches seek to generalize knowledge gained from a limited set of training data to novel situations, enabling the interpretation of new information without explicit prior exposure. To achieve effective zero-shot prediction, the model often utilizes a compositional representation of the output [25, 26]. Compositional representation refers to the ability to understand and generate novel combinations of previously learned concepts or features. By learning the underlying structure and relationships between different elements, the model can generalize its knowledge to new, unseen instances.
@@
Visual image reconstruction is another prominent example of zero-shot prediction in brain decoding. This task aims to recover perceived novel images that were not encountered during the training phase, effectively reconstructing visual experiences from brain activity patterns [27, 28]. As our perceptual visual experiences cannot be fully covered by limited brain data, reconstruction methods require strong generalizability. Miyawaki et al. [28] conducted a study demonstrating the reconstruction of perceived arbitrary $10\times 10$ binary-contrast images from brain activity. They built multiple modular decoders to predict the local contrasts of each location and combined their predictions. This approach leverages the compositional representation of the visual field, which is organized retinotopically in the early visual cortex. Incorporating cortical organization into the model’s architecture can significantly improve its ability to perform zero-shot prediction and reconstruct novel visual experiences from brain activity. Although the training stimuli were only $400$ random images, it was possible to reconstruct an arbitrary image from a set of possible $2^{100}$ instances, including geometric shapes such as crosses and alphabets. Similarly, Shen et al. [29] replaced local decoders with DNN feature decoders. Although their training stimuli were $1,200$ natural images, they demonstrated reconstructing novel images, including artificial images, which were not part of the training set. These successes suggest that the proposed reconstruction models capture rich and comprehensive information about the general aspects of the neural code, beyond merely the information defined by the training data [30]. Developing reliable reconstruction methods also enables further analysis of subjective visual experiences, such as visual imagery [29], attention [31], and illusion [32]. Enabling the decoding of novel brain states that were never encountered during the training phase can be a promising approach to neural mind-reading [2].
@@
Visual image reconstruction pipelines typically comprise three main components: translator, latent features, and generator (Fig 1). The translator converts brain activity patterns into a latent feature space, employing either linear regression [29, 33, 34, 35] or nonlinear transformation [36]. Latent features serve as surrogate representations of perceived visual images, evolving from primitive forms like local contrasts [28] to more sophisticated DNN features, such as intermediate outputs of recognition models [19, 29]. In our current work, we reframe this process as “translation” rather than “feature decoding,” a term we used in previous studies. This terminology acknowledges two important points: first, brain activity itself can be considered a latent representation of an image or perception of it, and second, it helps avoid potential ambiguity between image encoding/decoding and brain encoding/decoding processes. This new perspective conceptualizes the process as a translation between two latent spaces: from neural representation to machine representation.
@@
Fig 1: Visual image reconstruction pipeline. The first step involves translating brain activity patterns into machine/AI latent representations. While this process typically involves brain decoding of latent features using machine learning, here we term it “translation.” This terminology acknowledges that brain activity itself can be viewed as a latent representation of an image, framing the process as a translation from neural to machine latent representations. In the second step, a generator module takes these translated latent features as input and converts them into a visual image that corresponds to the content represented by the original brain activity.
@@
The generator visualizes translated latent features into images. Some studies have used pretraining image generative models for the generator module [34, 36, 35]. Image optimization can also be regarded as a generator [29]. End-to-end mapping from brain activity to images using DNNs can also be considered to contain these components implicitly or as a generator-only method [37, 38, 39, 40, 41, 42, 43].
@@
Recent advances in generative AI, particularly in text-to-image generation, have naturally given rise to expectations that these techniques could provide a valuable tool for visual image reconstruction by leveraging semantic representations. In addition, there has been a growing trend towards collecting neural datasets using a wide range of diverse visual and semantic content. This shift aims to capture a more comprehensive and ecologically valid representation of the human experience [13]. Researchers have started to collect large-scale fMRI datasets, such as the Natural Scene Dataset (NSD) [12] and the THINGS-fMRI dataset [14], which incorporate a broader range of brain data induced by diverse visual stimuli with text or category annotations. Notably, recent studies leveraging the latest generative AI techniques with semantic representation and large-scale datasets have reported realistic reconstructions from brain activity [44, 45, 46, 47, 48]. They utilized contrastive language-image pretraining (CLIP) features [22] as latent features and text-to-image diffusion models [23, 49] as generators.
@@
In the following, we begin with a case study that critically tests the methods proposed by Takagi and Nishimoto [44] and Ozcelik and VanRullen [45], which originally used the NSD. Our findings reveal several issues, including the failure to replicate the same level of reconstruction with another dataset specifically designed to avoid overlap between training and test sets [29], and the questionable post-hoc selection procedures of Takagi and Nishimoto that can produce convincing reconstructions even with random brain data. We identify the lack of diversity represented by the limited number of semantic clusters in the NSD dataset as a potential factor contributing to these issues. We also demonstrate the failure of zero-shot prediction in the latent feature space and the inability to recover a stimulus from its latent features, suggesting a misspecification of the model. Based on these findings, we conclude that the apparent realistic reconstructions are essentially a result of classification to clusters shared between training and test sets, combined with hallucinations by the generative model.
@@
In the formal analysis and simulation section, we aim to uncover the general factors underlying the issues and observations highlighted in the case study. We first describe the phenomenon of “output dimension collapse” in the translation from brain activity to the latent feature space. We show that the regression model trained on the clustered targets becomes specialized to the training examples, and its outputs collapse into a subspace of the training set. Our simulations with clustered data demonstrate that out-of-sample prediction can be achieved with an increasing number of training clusters, suggesting that zero-shot prediction is possible if stimulus diversity and a compositional representation are available. We also discuss the caveats associated with evaluating reconstructions using identification metrics alone and explore the preservation of image information at hierarchical layers of DNNs. Finally, we provide speculation on how we are fooled by the seemingly realistic reconstructions generated by AI models.
We conclude with recommendations for critically testing reconstruction models. Our findings highlight the importance of rigorous evaluation, the use of diverse datasets, and the need for careful specification and characterization of model components to advance the field of visual image reconstruction and develop truly generalizable prediction capabilities.
Results
Case study
@@ Results
We investigated two types of recent generative AI-based reconstruction methods, StableDiffusionReconstruction [44] and Brain-Diffuser [45], as well as their validation dataset, the Natural Scene Dataset (NSD) [12]. Both reconstruction methods utilize CLIP features [22] to effectively apply recent text-to-image diffusion models in visual image reconstruction analysis. CLIP text features are obtained from the average of five text annotations corresponding to the stimulus image. This text annotation information is only used during training to map the brain activity into the CLIP text features. In the test phase, they directly translate text features from brain activity during image perception. Hence, these two reconstruction methods will be referred to together as text-guided reconstruction methods.
@@
The StableDiffusionReconstruction method [44] uses the components of the Stable Diffusion model [50], the VAE features [51], and CLIP text features as latent features. Similarly, the Brain-Diffuser method [45] utilizes the component of another type of diffusion model [49], CLIP text features, CLIP vision features, and VDVAE features [52] as latent features. Both methods first generate initial images from translated VAE/VDVAE features. These initial images are then passed through the image-to-image pipeline of the diffusion model conditioned on the translated CLIP features, producing the final reconstructed images. They validated the reconstruction performance using the NSD dataset, preparing training and test data based on the data split provided by the NSD study. Thanks to the authors of these studies who made efforts to make their datasets and scripts publicly available, we were able to conduct our replication analysis effectively. We compared the reconstructed results of these two text-guided reconstruction methods with those from a previous image reconstruction method, iCNN [29]. For more information on datasets and reconstruction methods, see Material and Methods (“Datasets” or “Reconstruction methods”) or the original studies [12, 29, 44, 45].
@@
Observations: Failed replication and convincing reconstruction from random data

### v5

@@ Abstract
Advances in brain decoding, particularly in visual image reconstruction, have sparked discussions about the societal implications and ethical considerations of neurotechnology. As reconstruction methods aim to recover visual experiences from brain activity and achieve prediction beyond training samples (zero-shot prediction), it is crucial to assess their capabilities and limitations to inform public expectations and regulations. Our case study of recent text-guided reconstruction methods, which leverage a large-scale dataset (Natural Scene Dataset, NSD) and text-to-image diffusion models, reveals critical limitations in their generalizability, demonstrated by poor reconstructions on a different dataset. UMAP visualization of the text features from NSD images shows limited diversity with overlapping semantic and visual clusters between training and test sets. We identify that clustered training samples can lead to “output dimension collapse,” restricting predictable output feature dimensions. While diverse training data improves generalization over the entire feature space without requiring exponential scaling, text features alone prove insufficient for mapping to the visual space. Our findings suggest that the apparent realism in current text-guided reconstructions stems from a combination of classification into trained categories and inauthentic image generation (hallucination) through diffusion models, rather than genuine visual reconstruction. We argue that careful selection of datasets and target features, coupled with rigorous evaluation methods, is essential for achieving authentic visual image reconstruction. These insights underscore the importance of grounding interdisciplinary discussions in a thorough understanding of the technology’s current capabilities and limitations to ensure responsible development.
keywords:
Brain decoding , Visual image reconstruction , Naturalistic approach , NeuroAI
journal: Neural Networks
[…]
[inst1]organization=Graduate School of Informatics, Kyoto University,addressline=Yoshida-honmachi, Sakyo-ku, city=Kyoto, postcode=606-8501, country=Japan
[…]
[inst2]organization=Computational Neuroscience Laboratories, Advanced Telecommunications Research Institute International,addressline=Seika-cho, city=Sorakugun, postcode=619-0288, country=Japan
[…]
[inst3]organization=Institute for Quantum Life Science, National Institutes for Quantum Science and Technology,city=Chiba, postcode=263-8555, country=Japan
[…]
[inst4]organization=Guardian Robot Project, RIKEN,addressline=Seika-cho, city=Sorakugun, postcode=619-0288, country=Japan
1 Introduction
@@
Brain decoding has been widely used in the neuroscience field, revealing specific contents of the mind (Haxby et al., 2001; Kamitani and Tong, 2005; Soon et al., 2008; Horikawa et al., 2013). As brain decoding is sometimes referred to as “mind-reading” in popular media (Somers, 2021; Whang, 2023; Raasch, 2023), it has attracted significant attention beyond the scientific community due to its potential for real-world applications in medicine and industry. Such neurotechnology has also started to affect future ethical discussions and legal regulations (UNESCO, 2023). To prevent misleading public expectations and policies, scientists need to carefully assess the current status of brain decoding techniques and clarify the possibilities and limitations.
@@
One of the major challenges in brain decoding is the limited amount of brain data we can collect. The current brain measurement devices are costly, yielding far less brain data than the amounts typically used in image or text processing within the field of computer science and AI (Deng et al., 2009; Schuhmann et al., 2022). Although we have gradually increased the amount of brain data per subject (Van Essen et al., 2012; Allen et al., 2022; Naselaris et al., 2021; Hebart et al., 2023; Xu et al., 2024), it remains impractical to collect brain data covering the full range of cognitive states and perceptual experiences. The scarcity of brain data limits the applicability and scalability of classification-based decoding approaches, which are primarily developed in the early stage of this field. Such approaches can only decode information confined to the same stimuli or predefined categories used in the training phase, rendering them insufficient for uncovering the neural representation under general or natural conditions.
@@
To overcome this limitation, several decoding methods have been developed to enable the prediction of novel contents from brain activities that are not encountered during the training phase. Kay et al. (2008) proposed a general visual decoding approach via a statistical encoding model that predicted fMRI voxel values from image features. It successfully identified novel test images from a set of $1,000$ candidates. Mitchell et al. (2008) utilized co-occurrence rates of specific verb sets for nouns and built a computational model to predict fMRI voxel values while thinking about nouns in presented line drawing images. Their model demonstrated the ability to predict voxel values for novel nouns not seen during the training phase. Brouwer and Heeger (2009) constructed a color-tuning model and predicted brain activity while the subjects were presented with color stimuli. As their training stimuli covered most of the color space, their methods successfully identified novel colors not included in the training dataset. Horikawa and Kamitani (2017) utilized deep neural network (DNN) features to decode brain activity measured while subjects perceived natural images. They showed successful prediction of novel object categories not encountered during the training phase.
@@
In the field of machine learning, “zero-shot” prediction refers to the ability of a model to accurately predict or classify novel contents not encountered during the training phase (Larochelle et al., 2008; Palatucci et al., 2009). This ability has emerged in various applications across different domains, including image classification (Radford et al., 2021), image generation (Ramesh et al., 2021), and natural language processing (Brown et al., 2020). The concept of zero-shot prediction can be considered analogous to brain decoding techniques that aim to interpret brain activity patterns associated with previously unseen stimuli or experiences. Both approaches seek to generalize knowledge gained from a limited set of training data to novel situations, enabling the interpretation of new information without explicit prior exposure. To achieve effective zero-shot prediction, the model often utilizes a compositional representation of the output (Lake et al., 2017; Higgins et al., 2018). Compositional representation enables the understanding and generating of novel features through the combination of previously learned ones. By learning the underlying structure and relationships between different features, the model can generalize its knowledge to new, unseen instances.
@@
Visual image reconstruction is another prominent example of zero-shot prediction in brain decoding. This task aims to recover perceived novel images that were not encountered during the training phase, effectively reconstructing visual experiences from brain activity patterns (Stanley et al., 1999; Miyawaki et al., 2008). As our perceptual visual experiences cannot be fully covered by limited brain data, reconstruction methods require strong generalizability. Miyawaki et al. (2008) conducted a study demonstrating the reconstruction of perceived arbitrary $10\times 10$ binary-contrast images from brain activity. They built multiple modular decoders to predict the local contrasts of each location and combined their predictions. This approach leverages the compositional representation of the visual field, which is organized retinotopically in the early visual cortex. Incorporating cortical organization into the model’s architecture can improve its ability to perform zero-shot prediction and reconstruct novel visual experiences from brain activity. Although the training stimuli were only $400$ random images, it was possible to reconstruct an arbitrary image from a set of possible $2^{100}$ instances, including geometric shapes such as crosses and alphabets. Similarly, Shen et al. (2019b) replaced local decoders with DNN feature decoders. Although their training stimuli were $1,200$ natural images, they demonstrated reconstructing novel images, including artificial images, which were not part of the training set. These successes suggest that the proposed reconstruction models capture rich and comprehensive information about the general aspects of the neural representation, beyond merely the information defined by the training data (Kriegeskorte and Douglas, 2019). Developing reliable reconstruction methods also enables further analysis of subjective visual experiences, such as visual imagery (Shen et al., 2019b), attention (Horikawa and Kamitani, 2022), and illusion (Cheng et al., 2023). Decoding novel brain states that were never encountered during the training phase can be a promising approach to neural mind-reading (Kamitani and Tong, 2005).
@@
Visual image reconstruction pipelines typically comprise three main components: translator, latent features, and generator (Fig. 1). The translator converts brain activity patterns into a latent feature space, employing either linear regression (Shen et al., 2019b; Seeliger et al., 2018; Mozafari et al., 2020; Ozcelik et al., 2022) or nonlinear transformation (Qiao et al., 2020). Latent features serve as surrogate representations of perceived visual images, evolving from primitive forms like local contrasts (Miyawaki et al., 2008) to more sophisticated DNN features, such as intermediate outputs of recognition models (Horikawa and Kamitani, 2017; Shen et al., 2019b). In our current work, we reframe this process as “translation” rather than “feature decoding,” a term we used in previous studies. This terminology acknowledges two important points: first, brain activity itself can be considered a latent representation of an image or the perception formed by it and second, it helps avoid potential ambiguity between image encoding/decoding and brain encoding/decoding processes. This new perspective conceptualizes the process as a translation between two latent spaces: from neural representation to machine representation.
@@
Fig. 1: Visual image reconstruction pipeline. The first step involves translating brain activity patterns into machine/AI latent representations. While this process typically involves brain decoding of latent features using machine learning, here we term it “translation.” This terminology acknowledges that brain activity itself can be viewed as a latent representation of an image, framing the process as a translation from neural to machine latent representations. In the second step, a generator module takes these translated latent features as input and converts them into a visual image that corresponds to the content represented by the original brain activity.
@@
The generator visualizes translated latent features into images. Some studies have used pretraining image generative models for the generator module (Mozafari et al., 2020; Qiao et al., 2020; Ozcelik et al., 2022). Image optimization can also be regarded as a generator (Shen et al., 2019b). End-to-end mapping from brain activity to images using DNNs can also be considered to contain these components implicitly or as a generator-only method (Fujiwara et al., 2013; Shen et al., 2019a; Beliy et al., 2019; Ren et al., 2021; Gaziv et al., 2022; Lin et al., 2022; Chen et al., 2023).
@@
Recent advances in generative AI, particularly in text-to-image generation, have naturally given rise to expectations that these techniques could provide a valuable tool for visual image reconstruction by leveraging semantic representations. In addition, there has been a growing trend towards collecting neural datasets using a wide range of diverse visual and semantic content. This shift aims to capture a more comprehensive and ecologically valid representation of the human experience (Naselaris et al., 2021). Researchers have started to collect large-scale fMRI datasets, such as the Natural Scene Dataset (NSD; Allen et al., 2022) and the THINGS-fMRI dataset (Hebart et al., 2023), which include more than $10,000$ brain samples per subject. These datasets incorporate a broader range of brain data induced by diverse visual stimuli with text or category annotations. Importantly, recent studies have demonstrated that combining large-scale datasets with generative AI techniques can lead to more realistic reconstructions from brain activity (Takagi and Nishimoto, 2023a; Ozcelik and VanRullen, 2023; Scotti et al., 2023; Bai et al., 2024; Benchetrit et al., 2024; Scotti et al., 2024). These approaches commonly utilize linear regression models as translators, contrastive language-image pretraining (CLIP) text features (Radford et al., 2021) as part of the latent features, and text-to-image diffusion models (Ramesh et al., 2021; Xu et al., 2022) as generators. MindEye2 (Scotti et al., 2024) has recently shown improved reconstruction performance on the NSD, using a nonlinear translator, latent features of a CLIP’s image model, and a fine-tuned generator. This approach also includes a refinement step that enhances the realism of reconstructed images.
@@
In the following, we begin with a case study that critically tests text-guided reconstruction methods. We evaluate the approaches of Takagi and Nishimoto (2023a) and Ozcelik and VanRullen (2023), which were originally evaluated using the NSD (additionally, Scotti et al. (2024)’s method is examined). Our analysis reveals limitations in these methods. First, reconstruction quality substantially degrades when tested on a dataset specially designed to avoid object category overlaps between training and test sets (Shen et al., 2019b). Second, the post-hoc selection procedure used by Takagi and Nishimoto (2023a) can produce seemingly convincing reconstructions even from random brain data when applied to the NSD. Further investigation reveals limited semantic and visual diversity in the NSD stimulus set, with few distinct semantic clusters, potentially explaining these issues. We also demonstrate the failure of zero-shot prediction in the latent feature space and the inability to recover a stimulus from its latent features, suggesting fundamental constraints in the latent feature representation. These findings indicate that the apparent realism of reconstructions likely results from classification into clusters shared between training and test sets, combined with hallucinations by the generative model.
@@
In the formal analysis and simulation section, we investigate the general factors underlying the limitations observed in our case study. We introduce the phenomenon of “output dimension collapse” that occurs when translating brain activity into latent feature space. Our analysis shows that regression models trained on clustered targets become overly specialized to training examples, causing their outputs to collapse into a restricted subspace of the training set. Through systematic simulations with clustered data, we demonstrate that successful out-of-sample prediction requires the number of training clusters to scale linearly with feature dimensionality, suggesting that zero-shot prediction becomes feasible given sufficient stimulus diversity and compositional representations. We also discuss the caveats associated with evaluating reconstructions using identification metrics alone and explore the preservation of image information at hierarchical layers of DNNs. Finally, we provide general accounts on how we could be fooled by seemingly realistic reconstructions generated by AI models. Based on these analyses, we conclude with recommendations for developing more reliable reconstruction methods and establishing rigorous evaluation protocols.
2 Results
2.1 Case study
@@
We primarily investigated two recent generative AI-based reconstruction methods, StableDiffusionReconstruction (Takagi and Nishimoto, 2023a) and Brain-Diffuser (Ozcelik and VanRullen, 2023), as well as their validation dataset, the Natural Scene Dataset (NSD; Allen et al., 2022). We selected these two methods for three key reasons. First, these methods represent reconstruction approaches that have gained significant public attention by leveraging recent advances in generative AI. Second, both utilize the NSD, which currently serves as a widely adopted benchmark for predictive modeling in the field. Third, these methods employ straightforward linear translators through ridge regression, an approach that has become standard practice. Both reconstruction methods utilize CLIP features (Radford et al., 2021) to effectively apply recent text-to-image diffusion models in visual image reconstruction analysis. CLIP text features are obtained from the average of five text annotations corresponding to the stimulus image. This text annotation information is only used during training to map the brain activity into the CLIP text features. In the test phase, they directly translate CLIP text features from brain activity during image perception. Hereafter, these two reconstruction methods will be referred to together as text-guided reconstruction methods. We also replicated the MindEye2 reconstruction method (Scotti et al., 2024), a more recent study evaluated using the NSD. Unlike the other two methods, MindEye2 uses a variant of the CLIP model’s image embeddings as latent features, implements a nonlinear translator, and utilizes specially designed diffusion-based generators rather than relying on text features as direct decoding targets.
@@
The StableDiffusionReconstruction method (Takagi and Nishimoto, 2023a) uses components of the Stable Diffusion model (Rombach et al., 2022), the VAE features (Kingma and Welling, 2014), and CLIP text features as latent features. Similarly, the Brain-Diffuser method (Ozcelik and VanRullen, 2023) utilizes components of another type of diffusion model (Xu et al., 2022), CLIP text features, CLIP vision features, and VDVAE features (Child, 2021) as latent features. Both methods translate brain activity into their latent features using a linear translator, and initial images are first generated from translated VAE/VDVAE features. These initial images are then passed through the image-to-image pipeline of the diffusion model conditioned on the translated CLIP features, producing the final reconstructed images. They validated the reconstruction performance using the NSD dataset, preparing training and test data based on the data split provided by the NSD study. Thanks to the authors’ efforts in making the datasets and scripts publicly available, we were able to conduct our replication analysis effectively. We compared the reconstructed results of these two text-guided reconstruction methods (additionally, MindEye2) with those from a previous image reconstruction method, iCNN (Shen et al., 2019b). For more information on datasets and reconstruction methods, refer to Methods (“Datasets” and “Reconstruction methods”) or the original studies (Allen et al., 2022; Shen et al., 2019b; Takagi and Nishimoto, 2023a; Ozcelik and VanRullen, 2023).
@@
2.1.1 Observations: Failed replication and convincing reconstruction from random data